find an equation whose roots are 1 more tahn the roots of6x2+1x-1 =0
Answer by ankor@dixie-net.com(22740)

find an equation whose roots are 1 more than the roots of 6x^2 + 1x - 1 = 0
:
Factor 6x^2 + 1x - 1 = 0
(3x - 1)(2x + 1) = 0
:
3x = +1
x = 1/3
and
2x = -1
x = -1/2
:
Add 1 to each root:
x = 1/3 + 4/3
and
x = -1/2 + 1 = +1/2
:
3x = 4
and
2x = 1
:
FOIL: (3x -4)(2x - 1) = 0
6x^2 - 3x - 8x + 4
6x^2 - 11x + 4 = 0 is the equation